# CLASH: Collaborative Large-Small Hierarchical Framework for Continuous Vision-and-Language Navigation

### v1

@@ Abstract
Vision-and-Language Navigation (VLN) requires robots to follow natural language instructions and navigate complex environments without prior maps. While recent vision-language large models demonstrate strong reasoning abilities, they often underperform task-specific panoramic small models in VLN tasks. To address this, we propose CLASH (Collaborative Large-Small Hierarchy), a VLN-CE framework that integrates a reactive small-model planner (RSMP) with a reflective large-model reasoner (RLMR). RSMP adopts a causal-learning-based dual-branch architecture to enhance generalization, while RLMR leverages panoramic visual prompting with chain-of-thought reasoning to support interpretable spatial understanding and navigation. We further introduce an uncertainty-aware collaboration mechanism (UCM) that adaptively fuses decisions from both models. For obstacle avoidance, in simulation, we replace the rule-based controller with a fully learnable point-goal policy, and in real-world deployment, we design a LiDAR-based clustering module for generating navigable waypoints and pair it with an online SLAM-based local controller. CLASH achieves state-of-the-art (SoTA) results (ranking 1-st) on the VLN-CE leaderboard, significantly improving SR and SPL on the test-unseen set over the previous SoTA methods. Real-world experiments demonstrate CLASH’s strong robustness, validating its effectiveness in both simulation and deployment scenarios.
@@ I Introduction
Given a natural instruction such as “Go out of the current room and enter the kitchen on your right,” Vision-and-Language Navigation (VLN) [anderson2018vision] tasks an embodied agent with interpreting natural-language instructions to navigate through complex, real-world environments—without access to prior maps. Early VLN studies were conducted in discrete environments using the Matterport3D simulator [chang2017matterport3d], where agents navigated among pre-defined panoramic viewpoints. To enhance real-world relevance, research has since shifted to continuous settings (VLN-CE) [krantz_beyond_2020], where agents perceive egocentric views and execute low-level motor commands. Over years of development, two dominant paradigms have emerged in VLN-CE: (1) end-to-end models that directly map raw visual and textual inputs to low-level actions [krantz_beyond_2020, he2024mee, zhang2024navid, cheng2024navila, wei2025streamvln]; and (2) hierarchical “brain-body” frameworks, where a high-level planner predicts navigation waypoints, which are then executed by a low-level controller [anderson2021sim, hong2022bridging, He2024_IAHWP, wang2023dreamwalker]. Initially, both paradigms relied on compact, task-specific small models, often enhanced by domain-specific data augmentation [wang2023scaling, wang2023res, wang2023pasts, li2022envedit], improved multimodal semantic perception [wu2024tmm, tan2024self, wang2024seat, wang2023dual], and better encoding of historical trajectories [hong2021vln, chen2021history, chen2022think, an2022bevbert].
@@
Recently, multimodal large language models (MLLMs) [Qwen2.5-VL, wang2024qwen2, chen2024internvl], pretrained on large-scale image–text datasets, have exhibited strong cross-modal reasoning abilities. Their potential to generalize to embodied navigation has sparked growing interest in the VLN community. Existing MLLM-based VLN methods also fall into two main paradigms. End-to-end methods (e.g., NaVid [zhang2024navid], UniNaVid [zhang2024uni], NaVILA [cheng2024navila]) fine-tune MLLMs to map monocular RGB observations directly to actions, treating visual inputs as continuous video streams. In contrast, hierarchical methods (e.g., NavGPT [zhou2024navgpt], MapGPT [chen2024mapgpt], InstructNav [long2024instructnav], OpenVLN [qiao2024open]) employ MLLMs for high-level decision-making under zero- or few-shot settings.
Our empirical comparisons across these paradigms reveal two consistent trends. (1) Hierarchical “brain–body” frameworks with panoramic inputs consistently outperform fully end-to-end pipelines with monocular RGB inputs (e.g., g3D-LF [wang2025g3d] achieves 61% SR, about 7% higher than NaVILA [cheng2024navila]). (2) Within hierarchical frameworks, lightweight task-specific models still surpass fine-tuned MLLMs by a similar margin despite their smaller scale (e.g., MAGIC [wang2024magic] attains 79% SR vs. NavGPT2’s 72% SR [zhou2024navgpt2]).
Nevertheless, while small panoramic models perform well in simulation, they may overfit to specific scene styles, leading to degraded performance when transferred to more diverse environments [vlnpe]. Moreover, their opaque decision-making and limited interpretability hinder deployment in dynamic real-world scenarios. These observations raise a central question: How can we harness the complementary strengths of MLLMs and panoramic task-specific models to develop VLN agents that are both robust and generalizable?
@@
The critical question then becomes: how to effectively integrate RSMP and RLMR outputs for robust navigation? We address this through an uncertainty-aware collaboration mechanism (UCM) that adaptively fuses both models by estimating action confidence—prioritizing the small model when confident and the large model when uncertain. Unlike previous methods using cross-entropy [wang2024magic, qiao2024llm], learned weights [liu2021efficient, NEURIPS2020_aab08546, feng2022uln] or multiple plausible outcomes [gal2016dropout, wang2024lookahead, koh2021pathdreamer] for uncertainty estimation, we adopt conformal prediction [shafer2008tutorial], which provides statistically calibrated confidence measures with guaranteed coverage for more reliable decision-making. The advantages of using CP for uncertainty estimation are: (i) providing statistically reliable confidence measures that enable quantifiable control over decision reliability; (ii) not requiring model modification or additional training; and (iii) avoiding additional network prediction rounds, which would incur computational overhead.
Beyond high-level reasoning, low-level execution is equally critical for bridging simulation and real-world deployment. Prior hierarchical VLN methods [an2024etpnav, wang2025g3d, wang2023gridmm] rely on simulated panoramic depth for waypoint prediction and use rule-based controllers that rotate randomly to escape collisions. While functional in simulation, this strategy suffers from two major drawbacks: (i) unsafe behavior, as random rotations and movements may cause collisions with obstacles; (ii) impracticality for real-world deployment, since consumer-grade panoramic cameras do not support pixel-aligned depth sensing required for accurate waypoint prediction.
To address these challenges, we introduce a dual-level solution. For simulation, we replace rule-based control with a fully learnable point-goal navigation policy based on DDPPO [wijmansdd2019ddppo], which demonstrates robust obstacle avoidance and deadlock recovery. For real-world deployment, where synchronized depth is unavailable and existing approaches [shi2025smartway, li2025ground] require inefficient multi-turn camera capture that severely disrupts navigation fluency, we propose a LiDAR-based clustering method for waypoint generation. This training-free approach applies K-means clustering over LiDAR cost maps to extract traversable candidates. For local execution, we adopt an online SLAM-based controller that ensures robust localization and dynamic obstacle avoidance.
@@
4) Superior Performance and Real-World Validation. CLASH achieves state-of-the-art results on R2R-CE [krantz_beyond_2020] and REVERIE-CE [qi2020reverie], ranking 1st on the VLN-CE leaderboard, with extensive real-world experiments validating strong sim-to-real transferability.
@@ II Related Work
The VLN task [anderson2018vision, krantz_beyond_2020, qi2020reverie, vlnpe] requires a robot to follow language instructions to navigate real-world environments. Early works expanded from single-camera to panoramic observations [fried2018speaker], significantly improving performance. Subsequent efforts have explored diverse strategies, such as data augmentation [li2022envedit, wang2023pasts, source2025tan, wang2023scaling], sensor fusion [wu2024tmm, wang2024seat], and structured history modeling [yu2025moss, an2022bevbert]. The continuous setting (VLN-CE)[krantz_beyond_2020] requires agents to operate in a fine-grained 3D space using low-level actions. Within VLN-CE, research generally follows two main paradigms:
(a) Single-camera visual input: Agents use first-person RGB or RGB-D views. Early methods employed cross-modal architectures with LSTMs for temporal modeling [He2025_MLANet, he2023learning, li2022reve-ce]. Recent work uses MLLMs for commonsense reasoning and generalization. While zero-shot performance is limited [wang2024qwen2, long2024instructnav], fine-tuning with in-domain navigation data [zhang2024navid, cheng2024navila, zhang2024uni] significantly boosts action prediction, outperforming classical end-to-end monocular models.
(b) Panoramic visual input: Another approach uses panoramic observations for 360-degree perception [anderson2021sim, hong2022bridging, He2024_IAHWP, wang2023dreamwalker]. Compared to single-camera settings, panoramic inputs can naturally provide richer environmental context, enhancing spatial awareness and decision-making. Panoramic inputs are typically used in hierarchical frameworks with high-level waypoint prediction and low-level action execution [an2024etpnav, krantz2021waypoint, shi2025smartway]. Some methods also integrate semantic mapping for exploration and trajectory reasoning [wang2025g3d, wang2023gridmm, an2022bevbert, wang2024lookahead]. However, most panoramic methods rely on task-specific small models. Recent efforts incorporating MLLMs in discrete panoramic settings [zhou2024navgpt, zhou2024navgpt2, chen2024mapgpt] still underperform specialized models. Additionally, VLN-Copilot [qiao2024llm] uses MLLMs to generate auxiliary guidance cues concatenated with navigation instructions to aid smaller models, but this approach demands extensive training data and ultimately depends on small models for inference. Although several recent studies [qiao2024open, long2024instructnav, shi2025smartway] explore zero-shot panoramic navigation with MLLMs, their prompt designs yield limited gains.
@@
Reliable uncertainty estimation plays a crucial role in enabling robust and safe decision-making for embodied reasoning and navigation agents. Some methods employ entropy-based estimation, which computes uncertainty from prediction entropy or cross-entropy values [wang2024magic, qiao2024llm]. While these measures are simple and efficient, they often lack statistical calibration, leading to overconfident predictions under distribution shift. Other methods leverage auxiliary-head or evidential modeling to learn additional network parameters for predicting confidence scores [liu2021efficient, NEURIPS2020_aab08546, feng2022uln]. Although capable of capturing both aleatoric and epistemic uncertainty, these methods require model modification, retraining, and ground-truth labels for supervision. Additionally, some methods [gal2016dropout, wang2024lookahead, koh2021pathdreamer] generate multiple plausible outcomes to represent uncertainty, but at the cost of significant computational overhead.
In contrast, we adopt conformal prediction (CP) [shafer2008tutorial, huang2024conformal] for uncertainty estimation. Unlike heuristic or architecture-dependent methods, CP provides statistically calibrated confidence measures with formal coverage guarantees, enabling principled control over decision reliability. Recent studies have demonstrated its effectiveness in robotic planning and LLM-based reasoning [renrobots], where CP-guided confidence thresholds determine when an agent should trust its own plan or seek external guidance. Building upon this foundation, our approach leverages conformal prediction to achieve dynamic collaboration between the small and large models.
@@
Building on the remarkable success of large language models, recent efforts in Embodied AI have increasingly focused on developing vision-language foundation models to endow robots with more generalizable perception, reasoning, and decision-making capabilities. Broadly speaking, current approaches fall into three technical paradigms: The first aims to build fully end-to-end MLLM-based systems that directly map raw observations to action sequences for task execution [kimopenvla, zitkovich2023rt, zhou2025chatvla], striving for simplicity and task generalization. The second adopts a hierarchical “brain-body” architecture, where a high-level MLLM (a cerebrum-like “big brain”) interprets complex language commands and makes abstract decisions, which are then executed by a low-level controller (a cerebellum-like “little brain”) [driess2023palm, intelligence2025pi, huang2024rekep, belkhale2024rth]. The third follows a multi-agent collaborative design, where a complex task is decomposed into smaller subtasks, each handled by an expert agent. Their outputs are then coordinated to guide the final execution [li2023camel, guo2024embodied, feng2025multi].
@@ III Vision-Language Navigation Definition
Figure 2: Overview of the CLASH framework. CLASH integrates a reactive small model and a reflective vision-language large model through an uncertainty-aware collaboration mechanism, with tailored waypoint and point-goal prediction schemes for simulation and real-world deployment.
@@ IV-A Reactive Small-Model Planner (RSMP)
VLN is a sequential decision-making problem that can be formulated as a partially observable Markov decision process (POMDP). It is essential to effectively represent and align visual, linguistic, and memory features. To improve robustness beyond purely observational learning, we first incorporate causal learning [wang2024causal], which helps decouple spurious dependencies and enhance generalization. In addition, since existing hierarchical VLN-CE methods [an2024etpnav, wang2025g3d, wang2024lookahead] mainly fuse the global topological graph with the instruction and often overlook fine-grained local cues, we adopt a dual-graph structure following prior work in the discrete setting [chen2022think, wang2023scaling, wang2023dual].
@@ IV-A1 Causal Learning
Generalization to unseen environments is crucial for VLN agents. However, conventional training often leads models to rely on spurious correlations in observational data, which may limit robustness under distribution shifts. Causal learning [pearl2018book, wang2024causal] offers a principled framework for reasoning beyond such correlations. Specifically, traditional methods model the observational likelihood $P(\mathcal{Y}|\mathcal{X})$, where $\mathcal{X}$ denotes multi-modal inputs (vision, language, and history) and $\mathcal{Y}$ represents the action prediction. This formulation does not explicitly account for confounders—latent factors that jointly influence inputs and outputs and may introduce shortcut dependencies. Causal inference introduces the do-operator to model interventional reasoning and decouple such entangled effects. Following [wang2024causal], confounders are categorized into two types: observable confounders (e.g., room-type cues in vision, directional keywords in language) and unobservable confounders (e.g., stylistic variations, linguistic patterns, trajectory tendencies). To better capture their respective influences in VLN-CE, we incorporate back-door and front-door adjustment modules into the RSMP.
@@
where $x^{\prime}$ represents potential input samples from the entire training feature space, and $m$ denotes mediator features. This adjustment is approximated through cross-sampling from global feature dictionaries and in-sampling from current inputs, implemented efficiently via multi-head attention mechanisms. This process can be formulated as $FA(\mathbf{x},\mathbf{x}^{\prime})=\mathbb{E}_{x^{\prime}}[\mathbf{x}^{\prime}]+\mathbb{E}_{m|x}[\mathbf{m}].$ By integrating both BA and FA into the network architecture, the model learns more reliable feature representations across modalities, enhancing generalization to unseen environments [wang2024causal].

### v2

@@ Abstract
Vision-and-Language Navigation in Continuous Environments (VLN-CE) requires robots to follow natural language instructions and navigate complex environments without prior maps. While recent vision-language large models demonstrate strong reasoning abilities, they often underperform task-specific panoramic small models in VLN tasks. To address this, we propose CLASH (Collaborative Large-Small Hierarchy), a VLN-CE framework that integrates a Reactive Small Model Planner (RSMP) with a Reflective Large Model Reasoner (RLMR). RSMP adopts a causal-learning-based dual-branch architecture to enhance generalization, while RLMR leverages panoramic visual prompting with chain-of-thought reasoning to support interpretable spatial understanding and navigation. We further introduce an Uncertainty-aware Collaboration Mechanism (UCM) that adaptively fuses decisions from both models. For obstacle avoidance, we replace the rule-based controller with a fully learnable point-goal policy in simulation. For real-world deployment, we design a LiDAR-based clustering module for generating navigable waypoints and pair it with an online SLAM-based local controller. CLASH achieves state-of-the-art (SoTA) results (ranking 1-st) on the VLN-CE online leaderboard and relatively improves SR and SPL on the test set by 13.79% and 13.73% over the previous SoTA method. Real-world experiments further demonstrate CLASH’s strong robustness without any in-domain fine-tuning, validating its effectiveness in both simulation and deployment scenarios. The code is available at https://crystalsixone.github.io/vln-clash.github.io/.
@@ I Introduction
Given a natural instruction such as “Go out of the current room and enter the kitchen on your right,” Vision-and-Language Navigation (VLN) [5] tasks an embodied agent with interpreting natural-language instructions to navigate through complex, real-world environments—without access to prior maps. Early VLN studies were conducted in discrete environments using the Matterport3D simulator [8], where agents navigated among pre-defined panoramic viewpoints. To enhance real-world relevance, research has since shifted to continuous settings (VLN-CE) [38], where agents perceive egocentric views and execute low-level motor commands. Over years of development, two dominant paradigms have emerged in VLN-CE: (1) end-to-end models that directly map raw visual and textual inputs to low-level actions [38, 25, 77, 13, 71]; and (2) hierarchical “brain-body” frameworks, where a high-level planner predicts navigation waypoints, which are then executed by a low-level controller [4, 29, 28, 56]. Initially, both paradigms relied on compact, task-specific small models, often enhanced by domain-specific data augmentation [69, 57, 61, 41], improved multimodal semantic perception [73, 55, 62, 60], and better encoding of historical trajectories [30, 10, 11, 2].
@@
Recently, multimodal large language models (MLLMs) [6, 65, 12], pretrained on large-scale image–text datasets, have exhibited strong cross-modal reasoning abilities. Their potential to generalize to embodied navigation has sparked growing interest in the VLN community. Existing MLLM-based VLN methods also fall into two main paradigms. End-to-end methods (e.g., NaVid [77], UniNaVid [76], NaVILA [13]) fine-tune MLLMs to map monocular RGB observations directly to actions, treating visual inputs as continuous video streams. In contrast, hierarchical methods (e.g., NavGPT [79], MapGPT [9], InstructNav [45], OpenVLN [49]) employ MLLMs for high-level decision-making under zero- or few-shot settings.
Our empirical comparisons across these paradigms reveal two consistent trends. (1) Hierarchical “brain–body” frameworks with panoramic inputs consistently outperform fully end-to-end pipelines with monocular RGB inputs (e.g., g3D-LF [66] achieves 61% SR, about 7% higher than NaVILA [13]). (2) Within hierarchical frameworks, lightweight task-specific models still surpass fine-tuned MLLMs by a similar margin despite their smaller scale (e.g., MAGIC [59] attains 79% SR vs. NavGPT2’s 72% SR [78]).
Nevertheless, while small panoramic models perform well in simulation, they may overfit to specific scene styles, leading to degraded performance when transferred to more diverse environments [64]. Moreover, their opaque decision-making and limited interpretability hinder deployment in dynamic real-world scenarios. These observations raise a central question: How can we harness the complementary strengths of MLLMs and panoramic task-specific models to develop VLN agents that are both robust and generalizable?
@@
The critical question then becomes: how to effectively integrate RSMP and RLMR outputs for robust navigation? We address this through an uncertainty-aware collaboration mechanism (UCM) that adaptively fuses both models by estimating action confidence—prioritizing the small model when confident and the large model when uncertain. Unlike previous methods using cross-entropy [59, 48], learned weights [44, 1, 17] or multiple plausible outcomes [20, 67, 35] for uncertainty estimation, we adopt conformal prediction [52], which provides statistically calibrated confidence measures with guaranteed coverage for more reliable decision-making. The advantages of using CP for uncertainty estimation are: (i) providing statistically reliable confidence measures that enable quantifiable control over decision reliability; (ii) not requiring model modification or additional training; and (iii) avoiding additional network prediction rounds, which would incur computational overhead.
Beyond high-level reasoning, low-level execution is equally critical for bridging simulation and real-world deployment. Prior hierarchical VLN methods [3, 66, 68] rely on simulated panoramic depth for waypoint prediction and use rule-based controllers that rotate randomly to escape collisions. While functional in simulation, this strategy suffers from two major drawbacks: (i) unsafe behavior, as random rotations and movements may cause collisions with obstacles; (ii) impracticality for real-world deployment, since consumer-grade panoramic cameras do not support pixel-aligned depth sensing required for accurate waypoint prediction.
To address these challenges, we introduce a dual-level solution. For simulation, we replace rule-based control with a fully learnable point-goal navigation policy based on DDPPO [72], which demonstrates robust obstacle avoidance and deadlock recovery. For real-world deployment, where synchronized depth is unavailable and existing approaches [53, 43] require inefficient multi-turn camera capture that severely disrupts navigation fluency, we propose a LiDAR-based clustering method for waypoint generation. This training-free approach applies K-means clustering over LiDAR cost maps to extract traversable candidates. For local execution, we adopt an online SLAM-based controller that ensures robust localization and dynamic obstacle avoidance.
@@
4) Superior Performance and Real-World Validation. CLASH achieves state-of-the-art results on R2R-CE [38] and REVERIE-CE [47], ranking 1st on the VLN-CE leaderboard, with extensive real-world experiments validating strong sim-to-real transferability.
@@ II Related Work
The VLN task [5, 38, 47, 64] requires a robot to follow language instructions to navigate real-world environments. Early works expanded from single-camera to panoramic observations [19], significantly improving performance. Subsequent efforts have explored diverse strategies, such as data augmentation [41, 61, 54, 69], sensor fusion [73, 62], and structured history modeling [75, 2]. The continuous setting (VLN-CE)[38] requires agents to operate in a fine-grained 3D space using low-level actions. Within VLN-CE, research generally follows two main paradigms:
(a) Single-camera visual input: Agents use first-person RGB or RGB-D views. Early methods employed cross-modal architectures with LSTMs for temporal modeling [27, 26, 42]. Recent work uses MLLMs for commonsense reasoning and generalization. While zero-shot performance is limited [65, 45], fine-tuning with in-domain navigation data [77, 13, 76] significantly boosts action prediction, outperforming classical end-to-end monocular models.
(b) Panoramic visual input: Another approach uses panoramic observations for 360-degree perception [4, 29, 28, 56]. Compared to single-camera settings, panoramic inputs can naturally provide richer environmental context, enhancing spatial awareness and decision-making. Panoramic inputs are typically used in hierarchical frameworks with high-level waypoint prediction and low-level action execution [3, 36, 53]. Some methods also integrate semantic mapping for exploration and trajectory reasoning [66, 68, 2, 67]. However, most panoramic methods rely on task-specific small models. Recent efforts incorporating MLLMs in discrete panoramic settings [79, 78, 9] still underperform specialized models. Additionally, VLN-Copilot [48] uses MLLMs to generate auxiliary guidance cues concatenated with navigation instructions to aid smaller models, but this approach demands extensive training data and ultimately depends on small models for inference. Although several recent studies [49, 45, 53] explore zero-shot panoramic navigation with MLLMs, their prompt designs yield limited gains.
@@
Reliable uncertainty estimation plays a crucial role in enabling robust and safe decision-making for embodied reasoning and navigation agents. Some methods employ entropy-based estimation, which computes uncertainty from prediction entropy or cross-entropy values [59, 48]. While these measures are simple and efficient, they often lack statistical calibration, leading to overconfident predictions under distribution shift. Other methods leverage auxiliary-head or evidential modeling to learn additional network parameters for predicting confidence scores [44, 1, 17]. Although capable of capturing both aleatoric and epistemic uncertainty, these methods require model modification, retraining, and ground-truth labels for supervision. Additionally, some methods [20, 67, 35] generate multiple plausible outcomes to represent uncertainty, but at the cost of significant computational overhead.
In contrast, we adopt conformal prediction (CP) [52, 31] for uncertainty estimation. Unlike heuristic or architecture-dependent methods, CP provides statistically calibrated confidence measures with formal coverage guarantees, enabling principled control over decision reliability. Recent studies have demonstrated its effectiveness in robotic planning and LLM-based reasoning [51], where CP-guided confidence thresholds determine when an agent should trust its own plan or seek external guidance. Building upon this foundation, our approach leverages conformal prediction to achieve dynamic collaboration between the small and large models.
@@
Building on the remarkable success of large language models, recent efforts in Embodied AI have increasingly focused on developing vision-language foundation models to endow robots with more generalizable perception, reasoning, and decision-making capabilities. Broadly speaking, current approaches fall into three technical paradigms: The first aims to build fully end-to-end MLLM-based systems that directly map raw observations to action sequences for task execution [34, 81, 80], striving for simplicity and task generalization. The second adopts a hierarchical “brain-body” architecture, where a high-level MLLM (a cerebrum-like “big brain”) interprets complex language commands and makes abstract decisions, which are then executed by a low-level controller (a cerebellum-like “little brain”) [16, 33, 32, 7]. The third follows a multi-agent collaborative design, where a complex task is decomposed into smaller subtasks, each handled by an expert agent. Their outputs are then coordinated to guide the final execution [40, 23, 18].
@@ III Vision-Language Navigation Definition
Figure 2: Overview of the CLASH framework. First, the model employs (a) a navigable waypoint predictor to generate candidate waypoints in the surrounding environment. It then produces action predictions using (b) RSMP and (c) RLMR. When the estimated uncertainty is high, the (d) uncertainty-aware collaboration mechanism (UCM) is activated to fuse the decisions of the two models. Finally, low-level navigation is executed via (e) a Point-Goal Navigation module.
@@ IV-A Reactive Small-Model Planner (RSMP)
VLN is a sequential decision-making problem that can be formulated as a partially observable Markov decision process (POMDP). It is essential to effectively represent and align visual, linguistic, and memory features. To improve robustness beyond purely observational learning, we first incorporate causal learning [63], which helps decouple spurious dependencies and enhance generalization. In addition, since existing hierarchical VLN-CE methods [3, 66, 67] mainly fuse the global topological graph with the instruction and often overlook fine-grained local cues, we adopt a dual-graph structure following prior work in the discrete setting [11, 69, 60].
@@ IV-A1 Causal Learning
Generalization to unseen environments is crucial for VLN agents. However, conventional training often leads models to rely on spurious correlations in observational data, which may limit robustness under distribution shifts. Causal learning [46, 63] offers a principled framework for reasoning beyond such correlations. Specifically, traditional methods model the observational likelihood $P(\mathcal{Y}|\mathcal{X})$, where $\mathcal{X}$ denotes multi-modal inputs (vision, language, and history) and $\mathcal{Y}$ represents the action prediction. This formulation does not explicitly account for confounders—latent factors that jointly influence inputs and outputs and may introduce shortcut dependencies. Causal inference introduces the do-operator to model interventional reasoning and decouple such entangled effects. Following [63], confounders are categorized into two types: observable confounders (e.g., room-type cues in vision, directional keywords in language) and unobservable confounders (e.g., stylistic variations, linguistic patterns, trajectory tendencies). To better capture their respective influences in VLN-CE, we incorporate back-door and front-door adjustment modules into the RSMP.
@@
where $x^{\prime}$ represents potential input samples from the entire training feature space, and $m$ denotes mediator features. This adjustment is approximated through cross-sampling from global feature dictionaries and in-sampling from current inputs, implemented efficiently via multi-head attention mechanisms. This process can be formulated as $FA(\mathbf{x},\mathbf{x}^{\prime})=\mathbb{E}_{x^{\prime}}[\mathbf{x}^{\prime}]+\mathbb{E}_{m|x}[\mathbf{m}].$ By integrating both BA and FA into the network architecture, the model learns more reliable feature representations across modalities, enhancing generalization to unseen environments [63].